# Joyce structures from quadratic differentials on the sphere

### v1

@@ 9. Symmetries of $g$ and the Joyce structure
$E=\sum_{a=1}^{2n}z_{a}\frac{\partial}{\partial z_{a}}+E_{V}$ (9.33)
@@
$W(\sigma)$ $=\frac{W(Q_{1}(x))}{2y_{0}}dx-\frac{Q_{1}(x)W(y_{0})}{2Q_{0}(x)}dx$
$=\Bigg{(}-\frac{2}{2m_{\infty}-3}\frac{Q_{1}(x)}{2y_{0}}-\frac{2x}{2m_{\infty}-3}\frac{Q_{1}^{\prime}(x)}{2y_{0}}-\frac{x}{2m_{\infty}-3}\frac{Q_{1}(x)Q_{0}^{\prime}(x)}{2Q_{0}^{3/2}}\Bigg{)}dx$
$=-d\Bigg{(}x\frac{Q_{1}(x)}{2y_{0}}\Bigg{)}.$ (9.34)
@@
$E(\theta_{a})=\int_{\gamma_{a}}E(\sigma)=0$ (9.35)
@@
In the period coordinates $\{z_{a},\theta_{a}\}_{a=1}^{2n}$ the push-forward of the $2$-form $\Omega_{I}$ is given by
$i\Omega_{I}=\sum_{a=1}^{2n}\sum_{b=1}^{2n}\omega_{ab}dz^{a}\wedge d\theta^{b}.$ (9.36)
@@
$d\theta_{a}=d\bigg{(}\oint_{\gamma_{a}}\sigma\bigg{)}$ (9.37)
@@
The complex manifold $M=\operatorname{Quad}(\bar{m})$ carries a (meromorphic) Joyce structure compatible with the standard period structure.
@@
The condition (1) is just the statement the formula (8.8) holds and $\Omega_{-}$ is compatible with $\omega$. Now given an isomonodromic flow $L=U+V/\hbar$, from the definition (2.3) of the quaternionic structure we have (JiK)(V)=0(J-iK)(V)=0 and (JiK)(U)=2V(J-iK)(U)=2V. If we let $U$ be the unique isomonodromic flow that pushes down to the $z_{a}$ coordinate vector field (we can do this since the isomonodromic flows span a distribution transverse to the vertical bundle for all $\hbar\in\mathbb{C}^{*}$) then by Proposition 9.7
$\Omega\bigg{(}\hskip-2.84526ptU+\frac{V}{\hbar},\cdot\hskip-2.84526pt\bigg{)}=\sum_{b=1}^{2n}\bigg{(}\frac{\omega_{ab}dz^{b}}{\hbar^{2}}+\frac{\omega_{ab}dz^{a}(U)d\theta^{b}-\omega_{ab}d\theta^{a}\big{(}U+\frac{V}{\hbar}\big{)}dz^{b}}{\hbar}\bigg{)}+O(\hbar^{0})=0$ (9.38)
so that
$V=\frac{\partial}{\partial\theta^{a}}$ (9.39)
@@
as required for (2). For (3) we have from (2.3), I(U)=iUI(U)=-iU so that
$d\pi(U)=\frac{\partial}{\partial z^{a}}=iI(d\pi(U)).$ (9.41)
@@ 10. Four simple poles and Painlevé VI
$Q_{0}(x)=\frac{a(w-1)}{x}-\frac{aw}{x-1}+\frac{a}{x-w}$ (10.2)
@@
The earlier arguments of §4 through §9 need to be modified slightly to compute the Joyce structure. The map $\mu_{0}$ is defined by differentiating the tautological section with the Gauss-Manin connection as before. By looking at the order of the poles one may see that only residues at $x=w$ will contribute to $\omega$, the $2$-form induced by the intersection pairings on the elliptic curves. We have the Puiseux expansion
$y_{0}$ $=a^{1/2}(x-w)^{-1/2}+O((x-w)^{1/2}).$ (10.4)
@@
Given this potential there are two linearly independent isomonodromic flows for (1.1). The vector field
@@
Next, we consider the behaviour at $x=\infty$. Note that $Q$ has a zero of order 3 at infinity. Write $A(x)=A_{M_{\infty}}^{(\infty)}x^{M_{\infty}}+O(x^{M_{\infty}-1})$. The leading orders of the terms in (6.1) are at most the integers indicated below
@@
$+\frac{1}{\hbar A(w)}\Bigg{(}2pq(q-1)\frac{\partial}{\partial q}+\bigg{(}q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\bigg{)}\frac{\partial}{\partial v}\Bigg{)}$
@@
$g$ $=da\odot\Bigg{(}\frac{\kappa}{2pq(q-1)}dw-\frac{A(w)}{2pq(q-1)}dq\Bigg{)}-\bigg{(}\frac{\partial p}{\partial a}\bigg{)}^{-1}dw\odot dv$
$+$ $\bigg{(}\frac{\partial p}{\partial a}\bigg{)}^{-1}\Bigg{(}\frac{q(q-1)S^{\prime}(q)-v}{A(w)}-\frac{\kappa\big{(}q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\big{)}}{2pq(q-1)A(w)}\Bigg{)}dw\odot dw$
$+\bigg{(}\frac{\partial p}{\partial a}\bigg{)}^{-1}\Bigg{(}\frac{\big{(}q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\big{)}}{{2pq(q-1)}}\Bigg{)}dw\odot dq.$ (10.15)
@@
$+\frac{1}{2}\frac{q(q-1)(q-w)}{w^{2}(w-1)^{2}}\bigg{(}1-\frac{w}{q^{2}}+\frac{w-1}{(q-1)^{2}}\bigg{)}$ (10.17)
@@
and satisfies $g(E,E)=0$. Such metrics are known to ([15], Proposition 5.4) take a form generalising that of Sparling and Tod [25] in any Plebański coordinate system. The linearised data on $M$ of the Joyce metric $g^{J}$ and Joyce function $F$ (see §5 of [6] for definitions) vanish, for this $g$. The easiest way to see this is that the formula (73) in [6] for $F$ vanishes for the general form of the metric given in [15].
@@ 11. Conclusion and outlook
We have described a family of Joyce structures and provided a geometric description for their complex hyper-Kähler metrics: Each metric $g\in\Gamma(\odot^{2}T^{*}X)$, is described as the equivalent data of an $\mathcal{O}(2)$-valued $2$-form on twistor space naturally arising via the intersection pairings on the family of curves defined by $X$. By construction, in the coordinates $\{a_{i}^{(\alpha)}\hskip-2.84526pt,w_{\alpha},q_{I},v_{I}\}$ the components of the metric can be calculated via residue formulae which reduce to finite sums of Laurent coefficients. Accordingly, these hyper-Kähler metrics are amenable to explicit calculation, especially with the assistance of computer algebra.
@@
There are some aspects of the Joyce structures constructed here that warrant discussion. The hyper-Kähler structures each admit a projectable hyper-Lagrangian foliation: a rank-$2n$ distribution Lagrangian for the sphere of complex structures and pushing-down to a foliation of $M$ Lagrangian for the intersection pairing $\omega$. In a subsequent article we will explain that this foliation is related to the Hodge structure of the curves $\Sigma_{0}(p)$.
@@ Acknowledgements
The author would like to thank their PhD supervisor Maciej Dunajski for guidance and patience, and also Tom Bridgeland and Fabrizio Del Monte for helpful discussions. The author is supported by Cambridge Australia Scholarships and gratefully acknowledges the hospitality of the Simons Center for Geometry and Physics, Stony Brook University at which some of the research for this paper was performed, during the programme Geometric, Algebraic, and Physical structures around the moduli of Meromorphic Quadratic Differentials. This work has been partially supported by STFC consolidated grant ST/X000664/1.
Appendix A A. Explicit Isomonodromic Flows
Here we write out in detail the generators of the isomonodromic flows of the scalar ODE (1.1) with potential $Q(x)$ detailed in §5, in terms of parameters $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$.
There are $n$ linearly independent flows which in the body of the article we refer to as isopotential flows. They are somewhat “trivial” in that they preserve $Q(x)$. They are given by, for $1\leq i\leq m_{\alpha}$ where $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$
$L_{i}^{(\alpha)}\coloneqq U_{i}^{(\alpha)}+\frac{V_{i}^{(\alpha)}}{\hbar}$ (A.1)
@@
$U_{i}^{(\alpha)}\coloneqq\frac{\partial}{\partial a_{i}^{(\alpha)}}$ (A.2)
and the $\hbar^{-1}$ component is given by
$V_{i}^{(\alpha)}\coloneqq-\sum_{I=1}^{n}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}\frac{\partial}{\partial v_{I}}.$ (A.3)
Next, there are flows satisfying (6.1) with $A=(x-q_{I})^{-1}$ for I=1,,nI=1,...,n. We denote by $V_{I}$ the unique flow satisfying (6.1) for this $A$ and $\dot{a_{i}}^{(\alpha)}=0$ for $1\leq i\leq m_{\alpha}$ where $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$. Explicitly
$L_{I}=U_{I}+\frac{V_{I}}{\hbar}$ (A.4)
@@
$V_{I}=$ $-2p_{I}\frac{\partial}{\partial q_{I}}-\sum_{\begin{subarray}{c}K=1\end{subarray}}^{n}\Bigg{(}\sum_{i=0}^{2m_{\infty}-7}\frac{T_{i}^{(\infty)}(q_{K})q_{K}^{i}}{2p_{K}}+\sum_{\alpha=1}^{N}\sum_{i=1}^{2m_{\alpha}-1}\frac{T_{i}^{(\alpha)}(q_{K}-w_{\alpha})}{2p_{K}(q_{K}-w_{\alpha})^{i}}$
$+\sum_{\alpha=1}^{N}\sum_{j=1}^{2m_{\alpha}-1}\frac{ja_{j}^{(\alpha)}}{2p_{K}(q_{K}-w_{\alpha})^{j+2}}\Bigg{)}\frac{\partial}{\partial v_{K}}+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\Bigg{(}\frac{Q_{0}^{\prime}(q_{K})}{2p_{K}(q_{K}-q_{I})}-\frac{p_{K}}{(q_{K}-q_{I})^{2}}\Bigg{)}\frac{\partial}{\partial v_{K}}$
$+\Bigg{(}\frac{Q_{0}^{\prime}(q_{I})v_{I}}{p_{I}}-R^{\prime}(q_{I})+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\frac{p_{K}}{(q_{I}-q_{K})^{2}}\Bigg{)}\frac{\partial}{\partial v_{I}},$ (A.5)
@@
$U_{I}=$ $\sum_{i=m_{\infty}-3}^{2m_{\alpha}-7}T_{i}^{(\infty)}(q_{I})\frac{\partial}{\partial a_{i}^{(\infty)}}+\sum_{\alpha=1}^{n}\sum_{i=m_{\alpha}+1}^{2m_{\alpha}-1}T_{i}^{(\alpha)}(q_{I}-w_{\alpha})\frac{\partial}{\partial{a_{i}^{(\alpha)}}}+\sum_{\alpha=1}^{N}\frac{1}{q_{I}-w_{\alpha}}\frac{\partial}{\partial w_{\alpha}}$
$-2v_{I}\frac{\partial}{\partial q_{I}}-\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\frac{1}{q_{K}-q_{I}}\frac{\partial}{\partial q_{K}}+\Bigg{(}\sum_{K\neq I}^{n}\frac{3}{2(x-q_{K})^{3}}+\sum_{K\neq I}^{n}\frac{v_{K}}{(q_{I}-q_{K})^{2}}+S^{\prime}(q_{I})\Bigg{)}\frac{\partial}{\partial v_{I}}$
$+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\Bigg{(}\frac{3}{2(q_{K}-q_{I})^{3}}-\frac{v_{K}}{(q_{K}-q_{I})^{2}}\Bigg{)}\frac{\partial}{\partial v_{K}}.$ (A.6)
@@
$T_{i}^{(\alpha)}(x)$ $=\sum_{k=i}^{2m_{\alpha}-1}(k-2)(k-i+1)!\frac{a_{k}^{(\alpha)}}{x^{k-i+2}}$ (A.7)
$T_{i}^{(\infty)}(x)$ $=(2i-2m_{\infty}+7)x^{2m_{\infty}-i-7}+\sum_{k=i+2}^{2m_{\infty}-7}(2i-k+2)x^{k-i-2}a_{k}^{(\infty)}$ (A.8)
@@
Appendix B B. Generic non-degeneracy of the curve $\Sigma(\xi,\hbar)$
In this appendix we show that the algebraic curve $\Sigma(\xi,\hbar)$, defined by $y^{2}=Q(x)$ is non-degenerate for generic $(\xi,\hbar)\in X\times\mathbb{C}^{*}$.
Proposition B.1.
@@
$\mathcal{P}(x)=\hbar^{2}\prod_{I=1}^{n}(x-q_{I})^{2}\,\prod_{\alpha=1}^{N}(x-w_{\alpha})^{2m_{\alpha}-1}\,Q(x).$ (B.1)
@@
$\lim_{\hbar\to 0}\mathcal{P}(x)=\prod_{I=1}^{n}(x-q_{I})^{2}\,\prod_{\alpha=1}^{N}(x-w_{\alpha})^{2m_{\alpha}-1}\,Q_{0}(x)$ (B.2)
which is a polynomial with nN1n-N-1 zeroes of multiplicity one by the assumption that $Q_{0}(x)$ has distinct zeroes and $n$ zeroes of multiplicity $2$ at $x=q_{I}$. Since the position of distinct zeroes will vary holomorphically with $\hbar$ it suffices to show that for small $\hbar\neq 0$, $\mathcal{P}(x)$ has a pair of distinct zeroes in a neighbourhood of $x=q_{I}$. We may set $v_{I}=0$ for every $I$. Define
$\mathcal{F}_{I}(x)\coloneqq\prod_{\begin{subarray}{c}I=1\\ I\neq J\end{subarray}}^{n}(x-q_{J})^{2}\prod_{\alpha}^{N}(x-w_{\alpha})$ (B.3)

### v2

@@ 9. Symmetries of $g$ and the Joyce structure
$E=\sum_{a=1}^{2n}z^{a}\frac{\partial}{\partial z^{a}}+E_{V},$ (9.43)
@@
$E(\sigma)$ $=\frac{E(Q_{1}(x))}{2y_{0}}dx-\frac{Q_{1}(x)E(y_{0})}{2Q_{0}(x)}dx$
$=\Bigg(-\frac{2}{2m_{\infty}-3}\frac{Q_{1}(x)}{2y_{0}}-\frac{2x}{2m_{\infty}-3}\frac{Q_{1}^{\prime}(x)}{2y_{0}}+\frac{x}{2m_{\infty}-3}\frac{Q_{1}(x)Q_{0}^{\prime}(x)}{2Q_{0}^{3/2}}\Bigg)dx$
$=-\frac{2}{2m_{\infty}-3}d\Bigg(x\frac{Q_{1}(x)}{2y_{0}}\Bigg).$ (9.44)
@@
$E(\theta^{a})=-\int_{\gamma_{a}}E(\sigma)=0$ (9.45)
@@
In the period coordinates $\{z^{a},\theta^{a}\}_{a=1}^{2n}$ the push-forward of the $2$-form $\Omega_{I}$ is given by
$i\Omega_{I}=-\sum_{a=1}^{2n}\sum_{b=1}^{2n}\omega_{ab}dz^{a}\wedge d\theta^{b}.$ (9.46)
@@
$dz^{a}=\bigg(\oint_{\gamma_{a}}\psi\bigg),\quad d\theta^{a}=d\bigg(-\oint_{\gamma_{a}}\sigma\bigg)$ (9.47)
@@
The complex manifold $M=\operatorname{Quad}(\bar{m})$ carries a Joyce structure compatible with the standard period structure, which is possibly singular on the complement of an open dense subset of $TM$.
@@
The condition (1) is just the statement the formula (8.8) holds and $\Omega_{-}$ is compatible with $\omega$. Now, given an isomonodromic flow $L=U+V/\hbar$, from the definition (2.3) of the quaternionic structure we have (JiK)(V)=0(J-iK)(V)=0 and (JiK)(U)=2V(J-iK)(U)=2V. If we let $U_{a}$ be the unique isomonodromic flow that pushes down to the $z_{a}$ coordinate vector field (we can do this since the isomonodromic flows span a distribution transverse to the vertical bundle for all $\hbar\in\mathbb{C}^{*}$) then by Proposition 9.8
$\Omega\bigg(\hskip-2.84526ptU_{a}+\frac{V_{a}}{\hbar},\cdot\hskip-2.84526pt\bigg)=\sum_{b=1}^{2n}\bigg(\frac{\omega_{ab}dz^{b}}{\hbar^{2}}-\frac{\omega_{ab}d\theta^{b}-\omega_{cb}d\theta^{b}\big(U_{a}+\frac{V_{a}}{\hbar}\big)dz^{c}}{\hbar}\bigg)+O(\hbar^{0})=0$ (9.48)
so that from the $\hbar^{-2}$ term
$V_{a}=\frac{\partial}{\partial\theta^{a}}$ (9.49)
@@
as required for (2). For (3) we have from (2.3), $I(U_{a})=-iU_{a}$ so that
$d\pi(U_{a})=\frac{\partial}{\partial z^{a}}=i(d\pi(I(U_{a}))).$ (9.51)
@@ 10. Four simple poles and Painlevé VI
$Q_{0}(x)=\frac{a(w-1)}{x}-\frac{aw}{x-1}+\frac{a}{x-w}.$ (10.2)
@@
The earlier arguments of §4 through §9 need to be modified slightly to compute the Joyce structure. The map $\mu_{0}$ is defined by differentiating the tautological section with the Gauss-Manin connection as before. By looking at the order of the poles, one may see that only residues at $x=w$ will contribute to $\omega$, the $2$-form induced by the intersection pairings on the elliptic curves. We have the Puiseux expansion
$y_{0}$ $=a^{1/2}(x-w)^{-1/2}+O((x-w)^{1/2})$ (10.4)
@@
Given this potential, there are two linearly independent isomonodromic flows for (1.1). The vector field
@@
Next, we consider the behaviour at $x=\infty$. Note that $Q$ has a zero of order three at infinity. Write $A(x)=A_{M_{\infty}}^{(\infty)}x^{M_{\infty}}+O(x^{M_{\infty}-1})$. The leading orders of the terms in (6.1) are at most the integers indicated below
@@
$+\frac{1}{\hbar A(w)}\Bigg(2pq(q-1)\frac{\partial}{\partial q}+\bigg(q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\bigg)\frac{\partial}{\partial v}\Bigg),$
@@
$g$ $=da\Bigg(\frac{\kappa}{2pq(q-1)}dw-\frac{A(w)}{2pq(q-1)}dq\Bigg)-\bigg(\frac{\partial p}{\partial a}\bigg)^{-1}dw\,dv$
$+$ $\bigg(\frac{\partial p}{\partial a}\bigg)^{-1}\Bigg(\frac{q(q-1)S^{\prime}(q)-v}{A(w)}-\frac{\kappa\big(q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\big)}{2pq(q-1)A(w)}\Bigg)dw\,dw$
$+\bigg(\frac{\partial p}{\partial a}\bigg)^{-1}\Bigg(\frac{\big(q(q-1)R^{\prime}(q)-p-\frac{Q_{0}^{\prime}(q)\kappa}{2p}-A(w)\frac{\partial p}{\partial w}\big)}{{2pq(q-1)}}\Bigg)dw\,dq.$ (10.15)
@@
$+\frac{1}{2}\frac{q(q-1)(q-w)}{w^{2}(w-1)^{2}}\bigg(1-\frac{w}{q^{2}}+\frac{w-1}{(q-1)^{2}}\bigg).$ (10.17)
@@
and satisfies $g(E,E)=0$. Such metrics are known ([16], Proposition 5.4) to take a form generalising that of Sparling and Tod [29] in any Plebański coordinate system. The linearised data on $M$ of the Joyce metric $g^{J}$ and Joyce function $F$ (see §5 of [7] for definitions) vanish, for this $g$. The easiest way to see this is that the formula (73) in [7] for $F$ vanishes for the general form of the metric given in [16].
@@ 11. Conclusion and outlook
We have described a family of Joyce structures and provided a geometric description for their complex hyper-Kähler metrics: Each metric is described as the equivalent data of an $\mathcal{O}(2)$-valued $2$-form on twistor space naturally arising via the intersection pairings on the family of curves defined by $X$. By construction, in the coordinates $\{a_{i}^{(\alpha)}\hskip-2.84526pt,w_{\alpha},q_{I},v_{I}\}$ the components of the metric can be calculated via residue formulae which reduce to finite sums of Laurent coefficients. Accordingly, these hyper-Kähler metrics are amenable to explicit calculation, especially with the assistance of computer algebra.
@@
There are some aspects of the Joyce structures constructed here that warrant discussion. The hyper-Kähler structures each admit a projectable hyper-Lagrangian foliation: a rank-$2n$ distribution Lagrangian for the sphere of complex structures and pushing-down to a foliation of $M$, Lagrangian for the intersection pairing $\omega$. In a subsequent article we will explain that this foliation is related to the Hodge structure of the curves $\Sigma_{0}(p)$.
@@ Acknowledgements
The author would like to thank their PhD supervisor Maciej Dunajski for guidance and patience, Tom Bridgeland and Fabrizio Del Monte for helpful discussions, and the referees for their careful reading and comments, which have substantially improved the manuscript. The author is supported by Cambridge Australia Scholarships and gratefully acknowledges the hospitality of the Simons Center for Geometry and Physics, Stony Brook University at which some of the research for this paper was performed, during the programme Geometric, Algebraic, and Physical structures around the moduli of Meromorphic Quadratic Differentials. This work has been partially supported by STFC consolidated grant ST/X000664/1.
Appendix A A. Twistor distributions and quaternionic structures
We prove the following standard fact (usually phrased in terms of the Obata connection [28], Section 6):
Proposition A.1.
The distribution $L(\hbar)$ given by (2.2) is integrable for all $\hbar$ if and only if the complex structures $I,J,K$ defined by (2.3) are integrable.
Proof.
The forward implication is straightforward since we can identify the limits of $L(\hbar)$ as $\hbar\to 0$ and $\hbar\to\infty$ as the $\pm i$-eigenspaces of $I$, while $L(\pm i)$ gives the $\pm i$-eigenspaces of $J$ and so on. Conversely, suppose $I$ and $J$ are integrable. There are functions $\alpha_{ab}{}^{c},\beta_{ab}{}^{c},\gamma^{\pm}_{ab}{}^{c}$, a,b,c=1,,2na,b,c=1,...,2n satisfying
$[U_{a},U_{b}]$ $=\sum_{c=1}^{2n}\alpha_{ab}{}^{c}U_{c},\quad[V_{a},V_{b}]=\sum_{c=1}^{2n}\beta_{ab}{}^{c}V_{c},$ (A.1)
$[U_{a}\pm iV_{a},U_{b}\pm iV_{b}]$ $=\pm i([U_{a},V_{b}]+[V_{a},U_{b}])+\sum_{c=1}^{2n}\alpha_{ab}{}^{c}U_{c}-\sum_{c=1}^{2n}\beta_{ab}{}^{c}V_{c}$
[…]
Adding the $\pm$ equations from (A.2), and using linear independence of $\{U_{a},V_{a}\}_{a=1}^{2n}$:
[…]
Lastly, we use these to calculate
[…]
where
[…]
Appendix B B. Explicit Isomonodromic Flows
We provide formulae for the generators of the isomonodromic flows of the ODE (1.1) with potential $Q(x)$ detailed in §5, in terms of parameters $\{a_{i}^{(\alpha)},w_{\alpha},v_{I},q_{I}\}$.
There are $n$ linearly independent flows satisfying definition 7.1, we called isopotential flows. They are somewhat “trivial” in that they preserve $Q(x)$. They are given by, for $1\leq i\leq m_{\alpha}$ where $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$
$L_{i}^{(\alpha)}\coloneqq U_{i}^{(\alpha)}+\frac{V_{i}^{(\alpha)}}{\hbar},$ (B.1)
@@
$U_{i}^{(\alpha)}\coloneqq\frac{\partial}{\partial a_{i}^{(\alpha)}},$ (B.2)
while the $\hbar^{-1}$ component is given by
$V_{i}^{(\alpha)}\coloneqq-\sum_{I=1}^{n}\frac{\partial p_{I}}{\partial a_{i}^{(\alpha)}}\frac{\partial}{\partial v_{I}}.$ (B.3)
Next, to complete the basis of $2n$ flows we take a flow satisfying (6.1) with $A=(x-q_{I})^{-1}$ for I=1,,nI=1,...,n. We denote by $V_{I}$ the unique flow satisfying (6.1) for this choice of $A$ and $\dot{a_{i}}^{(\alpha)}=0$ for $1\leq i\leq m_{\alpha}$ where $\alpha=1,...,N$ and $0\leq i\leq m_{\infty}-4$ for $\alpha=\infty$. Explicitly
$L_{I}=U_{I}+\frac{V_{I}}{\hbar}$ (B.4)
@@
$V_{I}=$ $-2p_{I}\frac{\partial}{\partial q_{I}}-\sum_{\begin{subarray}{c}K=1\end{subarray}}^{n}\Bigg(\sum_{i=0}^{2m_{\infty}-7}\frac{T_{i}^{(\infty)}(q_{I})q_{K}^{i}}{2p_{K}}+\sum_{\alpha=1}^{N}\sum_{i=1}^{2m_{\alpha}-1}\frac{T_{i}^{(\alpha)}(q_{I}-w_{\alpha})}{2p_{K}(q_{K}-w_{\alpha})^{i}}$
$+\sum_{\alpha=1}^{N}\sum_{j=1}^{2m_{\alpha}-1}\frac{ja_{j}^{(\alpha)}}{2p_{K}(q_{I}-w_{\alpha})(q_{K}-w_{\alpha})^{j+1}}\Bigg)\frac{\partial}{\partial v_{K}}$ (B.5)
$+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\Bigg(\frac{Q_{0}^{\prime}(q_{K})}{2p_{K}(q_{K}-q_{I})}-\frac{p_{K}}{(q_{K}-q_{I})^{2}}\Bigg)\frac{\partial}{\partial v_{K}}$
$+\Bigg(\frac{Q_{0}^{\prime}(q_{I})v_{I}}{p_{I}}-R^{\prime}(q_{I})+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\frac{p_{K}}{(q_{K}-q_{I})^{2}}\Bigg)\frac{\partial}{\partial v_{I}},$ (B.6)
@@
$U_{I}=$ $\sum_{i=m_{\infty}-3}^{2m_{\infty}-7}T_{i}^{(\infty)}(q_{I})\frac{\partial}{\partial a_{i}^{(\infty)}}+\sum_{\alpha=1}^{N}\sum_{i=m_{\alpha}+1}^{2m_{\alpha}-1}T_{i}^{(\alpha)}(q_{I}-w_{\alpha})\frac{\partial}{\partial{a_{i}^{(\alpha)}}}+\sum_{\alpha=1}^{N}\frac{1}{q_{I}-w_{\alpha}}\frac{\partial}{\partial w_{\alpha}}$
$-2v_{I}\frac{\partial}{\partial q_{I}}-\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\frac{1}{q_{K}-q_{I}}\frac{\partial}{\partial q_{K}}+\Bigg(\sum_{K\neq I}^{n}\frac{v_{K}}{(q_{K}-q_{I})^{2}}-\sum_{K\neq I}^{n}\frac{3}{2(q_{K}-q_{I})^{3}}-S^{\prime}(q_{I})\Bigg)\frac{\partial}{\partial v_{I}}$
$+\sum_{\begin{subarray}{c}K=1\\ K\neq I\end{subarray}}^{n}\Bigg(\frac{3}{2(q_{K}-q_{I})^{3}}-\frac{v_{K}}{(q_{K}-q_{I})^{2}}\Bigg)\frac{\partial}{\partial v_{K}}.$ (B.7)
@@
$T_{i}^{(\alpha)}(x)$ $=\sum_{k=i}^{2m_{\alpha}-1}(2i-k-2)\frac{a_{k}^{(\alpha)}}{x^{k-i+2}},$ (B.8)
$T_{i}^{(\infty)}(x)$ $=(2i-2m_{\infty}+7)x^{2m_{\infty}-i-7}+\sum_{k=i+2}^{2m_{\infty}-7}(2i-k+2)x^{k-i-2}a_{k}^{(\infty)}$ (B.9)
@@
Appendix C C. Generic non-degeneracy of the curve $\Sigma(\xi,\hbar)$
In this appendix, we show that the algebraic curve $\Sigma(\xi,\hbar)$, defined by $y^{2}=Q(x)$ is non-degenerate for generic $(\xi,\hbar)\in X\times\mathbb{C}^{*}$.
Proposition C.1.
@@
$\mathcal{P}(x)=\hbar^{2}\prod_{I=1}^{n}(x-q_{I})^{2}\,\prod_{\alpha=1}^{N}(x-w_{\alpha})^{2m_{\alpha}-1}\,Q(x).$ (C.1)
@@
$\lim_{\hbar\to 0}\mathcal{P}(x)=\prod_{I=1}^{n}(x-q_{I})^{2}\,\prod_{\alpha=1}^{N}(x-w_{\alpha})^{2m_{\alpha}-1}\,Q_{0}(x),$ (C.2)
which is a polynomial with nN1n-N-1 zeroes of multiplicity one by the assumption that $Q_{0}(x)$ has distinct zeroes and $n$ zeroes of multiplicity $2$ at $x=q_{I}$. Since the position of distinct zeroes will (locally) vary holomorphically with $\hbar$, it suffices to show that for small $\hbar\neq 0$, $\mathcal{P}(x)$ has a pair of distinct zeroes in a neighbourhood of $x=q_{I}$. We may set $v_{I}=0$ for every $I$. Define
$\mathcal{F}_{I}(x)\coloneqq\prod_{\begin{subarray}{c}I=1\\ I\neq J\end{subarray}}^{n}(x-q_{J})^{2}\prod_{\alpha=1}^{N}(x-w_{\alpha})^{2m_{\alpha}-1}.$ (C.3)